OpenStudy (anonymous):

∭E xz dV,where E is the solid tetrahedron with vertices (0,0,0),(1,0,0),(1,1,0) and (1,0,1)

OpenStudy (loser66):

this is not exactly what you need, but I think it may help. It is from my cal2 course, you can take a look for fun.

OpenStudy (anonymous):

Thanks anyway Loser66. I'm having a little trouble finding the integration boundaries. Everything I try seems that the result isn't right. :\

OpenStudy (anonymous):

Anyone can confirme if the correct boudaries are: \[0 \le x \le 1\] \[0 \le y \le x\] \[0 \le z \le x-y\]